(D2+3D+2)y=cos2x
1
Expert's answer
2021-06-28T04:02:13-0400
"y''+3y'+2y=\cos2x"


Write the related homogeneous or complementary equation:


"y''+3y'+2y=0"


The general solution of a nonhomogeneous equation is the sum of the general solution "y_h(x)" of the related homogeneous equation and a particular solution "y_p(x)" of the nonhomogeneous equation:


"y(x)=y_h(x)+y_p(x)"

Consider a homogeneous equation 


"y''+3y'+2y=0"

Write the characteristic (auxiliary) equation:


"r^2+3r+2=0"

"(r+2)(r+1)=0"

"r_1=-2, r_2=-1"

The general solution of the homogeneous equation is


"y_h(x)=C_1e^{-2x}+C_2e^{-x}"


Let

"y_p=A\sin 2x+B\cos 2x"

Then


"y_p'=2A\cos 2x-2B\sin 2x"

"y_p''=-4A\sin2x-4B\cos 2x"

"+x(-A\sin x-B\cos x)"


Substitute


"-4A\sin2x-4B\cos 2x+6A\cos 2x-6B\sin 2x"

"+2A\sin 2x+2B\cos 2x=\cos 2x"

"-2A-6B=0"

"6A-2B=1"


"A=\dfrac{3}{20}, B=-\dfrac{1}{20}"



The general solution of a second order nonhomogeneous differential equation be


"y(x)=C_1e^{-2x}+C_2e^{-x}+\dfrac{3}{20}\sin 2x-\dfrac{1}{20}\cos 2x"